# An $\ell^{1}$-Plug-and-Play Approach for MPI Using a Zero Shot Denoiser with Evaluation on the 3D Open MPI Dataset

### v2

@@ Results
In Figure 5 a selected slice of the reconstructed shape phantom with different levels of preprocessing are displayed. We observe that the proposed ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP methods are rather stable with respect to different preprocessing choices. Further, the results suggest the possibility of employing the proposed methods with less preprocessed data. In the next experiment we go a step further, not relying on preprocessing nor on the SVD.
3.6 Experiment 4: Reconstruction With neither Preprocessing nor SVD
In our fourth and last experiment we study the proposed ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP algorithms on the OpenMPI data set with the minimum amount of preprocessing possible. In contrast to Experiment 2, we do not rely on any kind of SVD, neither for preprocessing, parameter selection, nor in the algorithmic part (Tikhonov type problem (9)). Employing methods not based on an SVD is important for future application scenarios since the computation of an SVD will become increasingly cumbersome with increasing problem size. The dimension of the system matrix in this scenario is $151230\times 6859$. In particular, without the SVD we cannot solve the Euler-Lagrange equation (9) directly, so we use the Conjugated Gradient algorithm with a tolerance of $10^{-12}$ and a number of maximum iterations of $10000$
Preprocessing
We do not perform any SNR thresholding, we do not whiten the data nor perform a rSVD. (However, we exclude the frequencies below 80 kHz.)
Parameters
We choose the parameters according to Section 2.4 and validate the number of iterations $n_{\mathrm{it}}$ and the parameter $\mu_{0}$. For the ZeroShot-PnP method the validated parameters are $n_{\mathrm{it}}=9$ and $\mu_{0}=10^{9}$ whereas for the ZeroShot-$\ell^{1}$-PnP $n_{\mathrm{it}}=48$ and $\mu_{0}=5\cdot 10^{10}$.
@@ 3.6 Experiment 4: Reconstruction With neither Preprocessing nor SVD
Figure 6: Selected slices of the reconstructions obtained using the proposed ZeroShot-$\ell^{1}$-PnP method (upper row) and with the ZeroShot-PnP method (lower row) with the validated parameters and no preprocessing of the data but the exclusion of the frequencies up to $80$ kHz.
@@
Table 6: Evaluation of the reconstructions using the validated parameters on the Open MPI dataset without preprocessing of the data.
Results
Selected slices for each phantom in the Open MPI Dataset are shown in Figure 6 and the quality evaluation with $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ is reported in Table 6. We observe that the result obtained in this experiments without preprocessing are comparable with the results obtained with the preprocessing and the knowledge of the SVD in Experiment 1 (cf. Table 2). The reconstruction time takes about 90 seconds, 30 seconds of which are taken to move the full matrix $A$ to the GPU and to form the matrix $A^{*}A$. Finally, the reconstruction with the full matrix with the ZeroShot-PnP (with and without the $\ell^{1}$-prior) approach here presented lasts about 1 minute including the computation of the quantity $A^{*}f$.
@@ 4.1 Discussion
We start out to discuss relations to the paper [4] which is to our knowledge the work closest to ours in the literature. The authors of [4] were the first to use a PnP method for MPI reconstruction. They train their denoiser on “MPI-friendly” MRI data using data of images of vessels, more precisely, they derived their training data from time-of-flight cerebral magnetic resonance angiograms (MRA) to closely mimic the vascular structures targeted in their MPI scans. In contrast to [4], we use a different approach for denoising. We employ a larger size neural network which has been trained on a larger size data set consisting of natural images. The employed denoiser is not learned or transfer learned on MPI or similar MRI data. Our results underpin that larger architectures with larger sets of training data can generalize well to situations where the data distribution is not exactly as in the training data, in particular for conservative tasks such as denoising. This is particularly appealing due to the very limited amount of (publicly) available MPI data. The authors of [4] provide a qualitative and quantitative evaluation on simulated data, and show the benefit of their PnP approach by qualitative evaluation on the resolution phantom of the Open MPI dataset. In contrast we provide a quantitative and qualitative evaluation of our ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP approach on the full 3D Open MPI dataset. Moreover, we use a different iterative scheme, namely the HQS, as a basis contrasting the ADMM type approach in [4]. Using HQS is appealing due to its simplicity and the ensuing relationship between the parameters and the Gaussian noise levels. Finally, we validate the parameters on a hybrid dataset. In addition, we discuss the relation with the DIP approach and the previous PP-MPI algorithm with trained denoiser [4]. The DIP approach stands out since, to our knowledge, it is the only machine learning approach for which a qualitative and quantitative evalution has been carried out on the Open MPI dataset [14]. The DIP approach is an unsupervised approach acting on a single image only. This is on the one hand appealing since other training data cannot influence the reconstruction and on the other hand no training data is needed which is particularly appealing in the context of MPI. Although not an unsupervised approach, the proposed method shares the property that we need no training data (due to using the denoiser in a zero shot fashion). However, the training data of the denoiser may influence the reconstruction. Moreover, the DIP approach suffers from high computational times whereas the ZeroShot-$\ell^{1}$-PnP can converge in a few seconds for preprocessed data and in about one minute for not preprocessed data. Finally, the benefit of using the zero-shot denoiser in the context of MPI is not only suggested by the quality of both the qualitative and quantitative results but also by its flexibility; in fact, in the previous PP-MPI approach the denoiser has been trained on MPI-friendly images of size $32\times 32$ pixels and with training data ranging between 0 and $1.5$. This means that, if the dimensionality of the MPI system-matrix inversion problem increases, it is reasonable to expect that another ad hoc denoiser shall be trained on a new dataset in order to be able to denoise more complicated images. The usage of a zero-shot denoiser pre-trained on a wide range of natural images might enable the employment of the suggested ZeroShot-PnP algorithm with little additional changes.
@@ 4.2 Conclusion
In this work, we have developed a PnP approach with $\ell^{1}$-prior for MPI reconstruction using a generic zero-shot denoiser. In particular, we employed a larger size DNN, namely the DRUnet architecture of [70] which has been trained on a larger size data set consisting of natural images. In order to apply the image denoiser on volumetric data, we employ slicing along the axes and averaging. We have quantitatively and qualitatively evaluated our approach on the full 3D Open MPI data set [39], which presently represents the only publicly available real MPI data set. To that end, we have employed suitable preprocessing, and we have compared our scheme quantitatively and qualitatively with the DIP approach of [14] as well as the Tikhonov regularization and the previous Plug-and-Play algorithm with trained denoiser [4]. Moreover, we have shown that the ZeroShot-$\ell^{1}$-PnP approach proposed works with parameters validated on a hybrid dataset and is able to perform fast reconstructions with different levels of preprocessing. In particular, we show that comparable results can be obtained without preprocessing and in particular without relying on SVD of the matrix.

### v3

@@ Results.
In Figure 6 a selected slice of the reconstructed shape phantom with different levels of preprocessing are displayed. We observe that the proposed ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP methods are rather stable with respect to different preprocessing choices. Further, the results suggest the possibility of employing the proposed methods with less preprocessed data. In the next experiment we go a step further, we consider only the basic preprocessing by backround correcting and suppressing excitation crosstalk and do not rely on the SVD.
3.6 Experiment 4: Reconstruction With neither SNR Thresholding, Whitening nor SVD.
In our fourth and last experiment we study the proposed ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP algorithms on the OpenMPI data set with the basic preprocessing consisting of background removal and suppression of excitation crosstalk. In contrast to Experiment 2, we do not rely on any kind of SVD, neither for preprocessing, parameter selection, nor in the algorithmic part (Tikhonov type problem (9)). Employing methods not based on an SVD is important for future application scenarios since the computation of an SVD will become incrementally cumbersome with increasing problem size. The dimension of the system matrix in this scenario is $151230\times 6859$. In particular, without the SVD we cannot solve the Euler-Lagrange equation (9) directly, so we use the Conjugated Gradient algorithm with a tolerance of $10^{-12}$ and a number of maximum iterations of $10000$
Preprocessing.
We do not perform any SNR thresholding, we do not whiten the data nor perform a rSVD. (We remind that the frequencies below 80 kHz are to be excluded because a hardware filtering is in place to suppress excitation crosstalk and reduces the SNR of frequencies below 80 kHz [54], cf. Section 2.2.)
Parameters.
We choose the parameters according to Section 2.4 and validate the number of iterations $n_{\mathrm{it}}$ and the parameter $\mu_{0}$. For the ZeroShot-PnP method the validated parameters are $n_{\mathrm{it}}=9$ and $\mu_{0}=10^{9}$ whereas for the ZeroShot-$\ell^{1}$-PnP $n_{\mathrm{it}}=48$ and $\mu_{0}=5\cdot 10^{10}$. The overview of the validated parameters can be found together with the reconstruction metrics in Table 7.
@@
Figure 7: Selected slices of the reconstructions obtained using the proposed ZeroShot-$\ell^{1}$-PnP method (upper row) and with the ZeroShot-PnP method (lower row) with the validated parameters and the basic preprocessing consisting of background correction and cutoff of the frequencies belo 80 kHz to suppress excitation crosstalk).
ZeroShot-PnP ZeroShot-$\ell^{1}$-PnP
Param. $\mu_{0}=10^{9}$ $\mu_{0}=5\cdot 10^{10}$
Iteration $n_{\mathrm{it}}=9$ $n_{\mathrm{it}}=48$
PSNR $33.87\pm 6.22$ $32.20\pm 5.08$
SSIM $0.831\pm 0.124$ $0.824\pm 0.135$
@@
Table 7: Validated parameters and evaluation of the reconstructions on the Open MPI dataset with only the cut-off of frequencies below 80 kHz as a preprocessing step.
Results.
Selected slices for each phantom in the Open MPI Dataset are shown in Figure 7 and the quality evaluation with $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ is reported in Table 7. We observe that the result obtained in this experiments with the minimal preprocessing (background correction and cutoff below 80 kHz) are comparable with the results obtained with the preprocessing and the knowledge of the SVD in Experiment 1 (cf. Table 2). The reconstruction time takes about 90 seconds, 30 seconds of which are taken to move the full matrix $A$ to the GPU and to form the matrix $A^{T}A$. Finally, the reconstruction with the full matrix with the ZeroShot-PnP (with and without the $\ell^{1}$-prior) approach here presented lasts about 1 minute including the computation of the quantity $A^{T}f$.
@@ 4.1 Discussion
We start out to discuss relations to the paper [4] which is to our knowledge the work closest to ours in the literature. The authors of [4] were the first to use a PnP method for MPI reconstruction. They train their denoiser on “MPI-friendly” MRI data using data of images of vessels, more precisely, they derived their training data from time-of-flight cerebral magnetic resonance angiograms (MRA) to closely mimic the vascular structures targeted in their MPI scans. In contrast to [4], we use a different approach for denoising. We employ a larger size neural network which has been trained on a larger size data set consisting of natural images. The employed denoiser is not learned or transfer learned on MPI or similar MRI data. This is particularly appealing due to the very limited amount of (publicly) available MPI data. The authors of [4] provide a qualitative and quantitative evaluation on simulated data, and show the benefit of their PnP approach by qualitative evaluation on the resolution phantom of the Open MPI dataset. In contrast we provide a quantitative and qualitative evaluation of the proposed ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP approach on the full 3D Open MPI dataset. Moreover, we use a different iterative scheme, namely the HQS, as a basis contrasting the ADMM type approach in [4]. The simplicity of the HQS splitting allows to easily formulate variants of it, including for example the $\ell^{1}$ norm, which can be substituted with no additional effort by another convex prior and its proximal map. An interesting observation from the results is that the additional $\ell^{1}$-prior does not play a prominent role in the reconstruction results, suggesting that the employment of the zero-shot denoiser and the parameter selection scheme are the main contributors to the reconstruction quality. Using HQS is appealing also due to relationship between the parameters and the Gaussian noise levels. Additionally, differently form the previous PP-MPI method, we employ a continuation strategy that allows, together with the interpretability of the parameters as Tikhonov regularization parameter and noise level, to produce an automatic parameter update at each iterations that feeds the neural network with the additional estimated noise level of the reconstruction at each iteration. Finally, we validate the parameters on a hybrid validation dataset and apply them to real data – in our case the Open MPI Dataset. In particular, for the validation dataset we consider generated phantoms whose maximum concentration are rescaled between 50 mmol and 150 mmol, which is the correct magnitude of the phantoms in the Open MPI dataset. We have observed that this validated dataset produced parameters that produced good quality results also for the resolution phantom, whose maximum concentration lies below 50 mmol. This suggests that a rough estimate of the concentration inside a target phantom seems to be enough to produce a reconstruction parameter choice using the same validation procedure on hybrid dataset but different rescalings.
In the previous PP-MPI approach the denoiser has been trained on MPI-friendly images of size $32\times 32$ pixels and with training data ranging between 0 and $1.5$. This means that, if the complexity of the MPI images to be reconstructed and the level of their details increases, it is reasonable to expect that another ad hoc denoiser shall be trained on a new dataset in order to be able to denoise more complicated images. The usage of a zero-shot denoiser pre-trained on a wide range of natural images might enable the employment of the suggested ZeroShot-PnP algorithm with little additional changes.
One interesting aspect of the proposed method is the automatic parameter selection at each iteration based on the estimation of the noise level and its input into the denoiser. The noise level estimation in this paper has been performed with the simple estimation formula in Eq.(13) which corresponds to the noise level of the iterate if the phantom is constant. This is of course not true for most phantoms and the formula in Equation (13) offers only an upper bound for the noise; in this direction, the employment of more sophisticated noise level estimation formulae appear to be effective.
Finally, we discuss relations with the DIP approach. The DIP approach stands out since, to our knowledge, it is the only machine learning approach for which a qualitative and quantitative evalution has been carried out on the Open MPI dataset [15]. The DIP approach is an unsupervised approach acting on a single image only. This is on the one hand appealing since other training data cannot influence the reconstruction and on the other hand no training data is needed which is particularly appealing in the context of MPI. Although not an unsupervised approach, the proposed method shares the property that we need no training data (due to using the denoiser in a zero shot fashion). However, the training data of the denoiser may influence the reconstruction. Moreover, the DIP approach suffers from high computational times whereas the ZeroShot-$\ell^{1}$-PnP can converge in a few seconds for preprocessed data and in about one minute for not preprocessed data. Finally, the benefit of using the zero-shot denoiser in the context of MPI is not only suggested by the quality of both the qualitative and quantitative results but also by its flexibility.
@@ 4.2 Conclusion
In this work, we have developed a PnP approach (with $\ell^{1}$-prior) for MPI reconstruction using a generic zero-shot denoiser. In particular, we employed a larger size DNN, namely the DRUnet architecture of [77] which has been trained on a larger size data set consisting of natural images. In order to apply the image denoiser on volumetric data, we employ slicing along the axes and averaging [4]. We have created a hybrid dataset that we use to estimate the reconstruction parameters; then, we applied the estimated parameters on the 3D Open MPI Dataset. We have quantitatively and qualitatively evaluated our approach on the full 3D Open MPI data set [43], which presently represents the only publicly available 3D Field-Free Point MPI data set. To that end, we have employed suitable preprocessing, and we have compared our scheme quantitatively and qualitatively with the DIP approach of [15] as well as the Tikhonov regularization, the ART method and the previous Plug-and-Play algorithm (PP-MPI) with trained denoiser [4]. We have shown that the proposed ZeroShot(-$\ell^{1}$)-PnP algorithmic scheme works well in combination with the ZeroShot-Denoiser [77] because of the ability of the ZeroShot-Denoiser to take into account noise level estimations of the reconstruction at each iteration of the algorithm; in relation to this, we have proposed an automatic parameter update strategy based on noise estimation for the ZeroShot-$\ell^{1}$-PnP algorithm. Moreover, we have shown that the proposed ZeroShot-($\ell^{1}$-)PnP approach is able to perform fast reconstructions with different levels of preprocessing. In particular, we show that comparable results can be obtained even only removing frequencies below 80 kHz and in particular without relying on the SVD of the system matrix.